# ToMoE: Converting Dense Large Language Models to Mixture-of-Experts through Dynamic Structural Pruning

### v1

@@ B.2 Efficient Implementation of $\mathcal{R}_{\mathbf{u}}$
$\mathbf{s}_{\mathbf{e}}=\text{ST-GS}(\text{Proj}_{\text{D}}^{\text{\tiny MLP}}(\mathbf{E})),$
@@ B.3 Equivalence of MoE and pseudo-MoE
One major challenge when training MoE models is maintaining an appropriate expert capacity, defined as the number of tokens each expert processes (Fedus et al., 2022). This is typically addressed using a load balancing loss. Without this loss, some experts may become overloaded while others remain underutilized, leading to bottlenecks where a few experts dominate the computation.
@@ Appendix C More Implementation Details
For the Alpaca dataset  [^1] https://huggingface.co/datasets/tatsu-lab/alpaca, we use the ‘text’ column within the dataset which combines the columns of ‘instruction’ and ‘output’. For the Code Alpaca dataset  [^2] https://github.com/sahil280114/codealpaca, we combine the ‘instruction’, ‘input’, and ‘output’ columns as one training sample.
@@ Appendix D More Experimental Results
In Tab. 6 and Tab. 7, we present the zero-shot performance of various methods on LLaMA-2 13B, Qwen-2.5 14B, and Phi-2 models. From Tab. 6, it is evident that ToMoE consistently outperforms other methods. Compared to 7B or 8B models, the performance gap between our method and other approaches is smaller, which also holds for the differences between baseline methods. This is likely due to the larger model sizes.

### v2

@@ B.2 Efficient Implementation of $\mathcal{R}_{\mathbf{u}}$
$\mathbf{s}_{\mathbf{e}}=\text{ST-GSig}(\text{Proj}_{\text{D}}^{\text{\tiny MLP}}(\mathbf{E})),$
@@
Table 13: Ablation study on design choices of ToMoE and the impact of temperature $\tau$ on performance.
Settings ARC-e ARC-c PIQA WinoG. HellaS. Avg
Local Emb 55.22 32.94 66.32 53.12 57.85 53.09
Head Pruning 45.12 25.85 64.82 49.49 39.65 44.99
w/o VO Routing 55.51 32.34 70.40 57.14 59.12 54.90
$\tau=0.3$ 57.28 33.53 70.35 57.54 60.22 55.78
$\tau=0.4$ 56.65 33.87 71.00 58.56 60.26 56.07
$\tau=0.5$ 55.89 33.36 71.00 56.75 59.31 55.16
ToMoE 56.65 33.87 71.00 58.56 60.26 56.07
@@ B.3 Equivalence of MoE and pseudo-MoE
One major challenge when training MoE models is maintaining an appropriate expert capacity, defined as the number of tokens each expert processes (fedus2022switch). This is typically addressed using a load balancing loss. Without this loss, some experts may become overloaded while others remain underutilized, leading to bottlenecks where a few experts dominate the computation.
@@ Appendix C More Implementation Details
For the Alpaca dataset  [^1] https://huggingface.co/datasets/tatsu-lab/alpaca, we use the ‘text’ column within the dataset, which combines the columns of ‘instruction’ and ‘output’. For the Code Alpaca dataset  [^2] https://github.com/sahil280114/codealpaca, we combine the ‘instruction’, ‘input’, and ‘output’ columns as one training sample.
@@ Appendix D More Experimental Results
Model Active Parameters Batch Size 512 Batch Size 1024 Batch Size 1536
LLaMA-2 7B 100% 1781.76 tokens/s 1805.80 tokens/s 1858.45 tokens/s
LLaMA-MoE E8A2 (zhu2024llama) 50% 2685.35 tokens/s 2796.07 tokens/s 2887.98 tokens/s
ToMoE (ours) 2360.38 tokens/s 2702.12 tokens/s 2919.45 tokens/s
Table 15: Inference throughput (tokens per second) under different mini-batchsizes.
In Tab. 9 and Tab. 10, we present the zero-shot performance of various methods on LLaMA-2 13B, Qwen-2.5 14B, and Phi-2 models. From Tab. 9, it is evident that ToMoE consistently outperforms other methods. Compared to 7B or 8B models, the performance gap between our method and other approaches is smaller, which also holds for the differences between baseline methods. This is likely due to the larger model sizes. Table 12 presents the zero-shot performance of the LLaMA-2 7B model across more baselines and active parameters. As shown in the table, ToMoE consistently achieves significantly better performance than all competing methods.
@@
Figure 10: MMLU accuracy vs. active parameters.
In Tab. 15, we report the inference throughput (measured in tokens per second) of different models under varying batch sizes. Compared to the dense LLaMA-2 7B baseline, both converted MoE models achieve higher throughput due to reduced active parameters. Our resulting model has a similar throughput to LLaMA-MoE when the batch size is large enough.
In Fig. 10, we present the accuracy–parameter trade-off on the more challenging MMLU dataset with the LLaMA-3 8B model. The results show that our method can still provide meaningful results when activating only 50% of the model parameters.
@@
In Tab. 13, we present an ablation study to analyze several design choices in ToMoE, focusing on architectural components and the effect of the temperature parameter $\tau$. Although ToMoE applies contextual sparsity to the value and output (VO) projections within the attention layer, their contribution to overall performance is relatively minor due to the small head dimension (128, in the case of LLaMA-2 7B). To validate this, we disable dynamic attention sparsity and instead apply only static structural pruning to the attention layer (denoted as "w/o VO routing"). This leads to only a modest performance drop of about 1% using 50% active parameters.
We also examine the role of global expert embeddings with GRU in conveying cross-layer architectural information. Specifically, we compare the default global expert embedding with a local-only variant ("local emb"), where expert embeddings are used only in MLP layers and removed from attention layers. Results show a slight decrease in performance, suggesting that global expert embeddings contribute to better coordination across layers.
Additionally, we evaluate the sensitivity of ToMoE to the temperature $\tau$ in the routing mechanism. The results with $\tau\in\{0.3,0.4,0.5\}$ show that performance remains relatively stable, indicating robustness to the choice of $\tau$ within a reasonable range.
Finally, we explore head pruning in the early stages of ToMoE development. However, this approach yielded significantly lower performance. This may be due to the distortion of attention feature maps when heads are removed, which makes it more difficult to train effective MLP experts.
@@ Appendix E Visualization of Experts
Appendix F Design Choice for the MHA
Ideally, to achieve maximum flexibility, one might consider applying dynamic pruning to all projection matrices in the MHA layer, including the query ($W_{Q}$), key ($W_{K}$), value ($W_{V}$), and output ($W_{O}$) matrices. However, there is a fundamental limitation when attempting to apply dynamic pruning along the head dimension for the query and key matrices.
Suppose we generate pruning masks $S_{t}\in\{0,1\}^{d}$ at each time step $t$ based on the input $X_{t}\in\mathbb{R}^{1\times d}$, and consider two distinct time steps, $a$ and $b$. For the $i$-th attention head, the attention score between queries and keys is influenced by the pruning masks. Specifically, the effective attention score between the $a$-th query and the $a$-th key is given by:
$e(X_{a}W_{Q,i})S_{a}S_{a}^{\top}e(X_{a}W_{K,i})^{\top},$
while the attention score between the $a$-th query and the $b$-th key is:
$e(X_{a}W_{Q,i})S_{a}S_{b}^{\top}e(X_{b}W_{K,i})^{\top}.$
The effective width—that is, the dimensionality over which attention is computed—between the $a$-th query and the $a$-th key is $\|S_{a}S_{a}^{\top}\|_{0}=\sum S_{a}=K$, assuming the mask has exactly $K$ active elements. However, for cross-position pairs like $(a,b)$, the effective width becomes $\|S_{a}S_{b}^{\top}\|_{0}=\|S_{a}\odot S_{b}\|_{0}\leq\min(\sum S_{a},\sum S_{b})=K$. The equality holds only when $S_{a}=S_{b}$, which generally does not hold for arbitrary $a\neq b$.
This observation implies that dynamically pruned query and key matrices fail to fully utilize the allocated capacity $K$ unless the pruning masks are identical across all positions. Moreover, the variability of the effective width across different query-key pairs introduces instability and inconsistent capacity utilization, making this approach less favorable compared to static pruning for the query and key matrices.
In contrast, dynamic pruning does not encounter this issue when applied to the value and output matrices, as these are not involved in pairwise comparisons like the query-key dot products. Therefore, we adopt dynamic pruning only for the value and output projections, while keeping the query and key projections pruned statically to maintain stable and full-capacity attention computation.